Franklin Manufacturing Company was started on January 1, 2018, when it acquired $90,000 cash by issuing common stock. Franklin immediately purchased office furniture and manufacturing equipment costing $9,800 and $31,400, respectively. The office furniture had an 8-year useful life and a zero salvage value. The manufacturing equipment had a $3,400 salvage value and an expected useful life of four years. The company paid $11,700 for salaries of administrative personnel and $15,900 for wages to production personnel. Finally, the company paid $11,540 for raw materials that were used to make inventory. All inventory was started and completed during the year. Franklin completed production on 4,200 unitsof product and sold 3,250 units at a price of $16 each in 2018. (Assume that all transactions are cash transactions and that product costs are computed in accordance with GAAP) Required a. Determine the total product cost and the average cost per unit of the inventory produced in 2018. (Round Average cost per unit to 2 decimal places.) b. Determine the amount of cost of goods sold that would appear on the 2018 income statement. (Do not round intermediate calculations.) the amount of the ending inventory balance that would appear on the December 31, 2018, balance sheet. (Do not round intermediate calculations.) d. Determine the amount of net income that would appear on the 2018 income statement. e. Determine the amount of retained earnings that would appear on the December 31, 2018, balance sheet f. Determine the amount of total assets that would appear on the December 31, 2018, balance sheet Average cost per unit b. Cost of good sold
Required a. Determine the total product cost and the average cost per unit of the inventory produced in 2018. (Round Average cost per unit to 2 decimal places.) b. Determine the amount of cost of goods sold that would appear on the 2018 income statement. (Do not round intermediate calculations.) c. Determine the amount of the ending inventory balance that would appear on the December 31, 2018, balance sheet. (Do not round intermediate calculations.) d. Determine the amount of net income that would appear on the 2018 income statement. e. Determine the amount of retained earnings that would appear on the December 31, 2018, balance sheet f. Determine the amount of total assets that would appear on the December 31, 2018, balance sheet. cost inventory d. Net
Answer #1
Annual Depreciation of office furnitures $1,225 (98000/8)
Annual Depreciation of office Equipment $        7,000 (31400-3400)/4
JOURNAL ENTRY
Account Title Debit Credit
Cash $90,000
Common Stock $90,000
Office Furniture $9,800
Cash $9,800
Equipment $31,400
Cash $31,400
Administration Salaries $11,700
$11,700
Production Wages $15,900
$15,900
Direct Material $11,540
Cash $11,540
Depreciation expense-Furniture $1,225
Accumulated Depreciation-Furniture $1,225
Depreciation expense-Equipment $        7,000
Accumulated Depreciation-Equipment $    7,000
Cash $52,000 (3250*16)
Sales Revenue $52,000
TOTAL PRODUCT COST
A Direct Material $11,540
B Direct Labor $15,900
C Manufacturing Overhead $7,000
D=A+B=C Total Product Cost $34,440
E Units Produced 4200
F=D/E Average Cost per unit $8.20
G Units Sold 3250
H=F*H Cost of goods sold $26,650
I=D-H Ending Inventory $7,790
NET INCOME
J Sales Revenue $52,000
K Cost of goods sold $26,650
L=J-K Gross Profit $25,350
M Administrative Salaries $11,700
N Depreciation-Furniture $1,225
P=L-N-M Net Income $12,425
a Total Product Cost $34,440
b Average Cost per unit $8.20
c Cost of goods sold $26,650
d Ending Inventory $7,790
e Net Income $12,425